# Powerformer: A Section-adaptive Transformer for Power Flow Adjustment

### v2

@@ 4.3 Comparison with Different Adjustment Baselines
To demonstrate the superiority of our method in efficiently and effectively adjusting transmission section power flow, we thoroughly analyze and compare our method with various popular approaches in deep reinforcement learning (DRL) framework, as well as a well-established traditional optimal power flow (OPF) adjustment method. Moreover, in order to thoroughly validate the effectiveness of our Powerformer, we design the additional 4-section power flow adjustment tasks, which aims to deal with the top four ranked sections in terms of transmission capacity among a total of ten sections.
@@
Through analyzing the learning curve in Fig. 4, our proposed method has successfully achieved the best test success rate among all considered architectures on three different systems. Although DQN and its variants require fewer timesteps to achieve reasonable results, Dueling DQN stands out with the highest achieved test success. This also provides advantageous conditions for our method to obtain the best results. We can analyze and compare the superiority of our method based on the difficulty levels of different tasks, especially in the more challenging task of 10-section adjustment (i.e., Fig. 4d4e and 4f), where the effectiveness of our method is highlighted. Furthermore, from the perspective of different system cases, our method performs even better on large-scale datasets. This is because our method is able to learn better representations of sectional power system states, thus avoiding the coupling of state features. We can comparatively analyze the learning curve on the IEEE 118-bus and European 9241-bus systems, and the significant differences in the effectiveness of our method can be directly observed because when the system case is simple, the task is also relatively simple, hence the superiority is not so apparent.
@@
From a quantifiable perspective of comparison adjustment algorithms, we leverage the comparative results of success rates and economic costs pertaining to 4-section and 10-section adjustment tasks, as evidenced in Table 3, to serve as prime examples for our comprehensive analysis. Specifically, we further analyze the results by examining different power system cases. For the IEEE 118-bus system, Powerformer outperforms all other methods with a success rate of 99.67%, significantly surpassing Dueling DQN (91.17%) on the 4-section task. As complexity increases in the 10-section task, Powerformer maintains a success rate of 98.19%, outperforming A2C (80.14%) by a substantial margin. For the China 300-bus system, Powerformer continues to exhibit a satisfying level of performance in comparison to the baseline approaches, accomplishing an impressively high test success rate of $99.73\%$ and $97.36\%$ on the 4-section and 10-section tasks, respectively. For the European-9241 system, our method consistently demonstrates exceptional performance, achieving an impressive success rate of 68.97% and 68.25% when considering the test success rates of 4-section and 10-section tasks. In contrast, alternative reinforcement learning methods fail to surpass the 35% threshold for the complex 10-section task. Additionally, our method consistently demonstrates its exceptional ability to achieve relatively low economic costs across all scenarios.
Figure 5: The inference time comparison of total 173 scenarios on the IEEE 118-bus system.
Table 4: The inference time of our method and OPF method on three different power systems. 118-bus System 300-bus System 9241-bus System OPF 44.822 $\pm$ 38.459 47.668 $\pm$ 28.230 1912.608 $\pm$ 1636.746 Powerformer (Ours) 0.078 $\pm$ 0.151 0.121 $\pm$ 0.176 0.585 $\pm$ 0.602
Figure 5: The inference time comparison of total 173 scenarios on the IEEE 118-bus system.
4.4 Inference Time Analysis
In real-time applications, such as the European 9241-bus system, inference time plays a critical role, particularly in large-scale systems. In such scenarios, power flow adjustment becomes more complex due to the numerous variations that need careful consideration and strategic decision-making. To better understand the superiority of our method in terms of inference speed, we provide Fig. 5 and Table 4 to demonstrate the advantages of our approach when compared to traditional OPF method used in practical scenarios. As shown in Fig. 5, the experimental results demonstrate that, when compared to the traditional OPF method, the time required to successfully achieve the predefined range for each individual test is significantly reduced across a comprehensive set of 173 test scenarios in the IEEE 118-bus system. Moreover, it is notable that our method exhibits a noteworthy reduction in inference time across three different power systems, as evidenced by the compelling numerical statistics presented in Table 4. As a result, our proposed method is remarkably more time efficient than the traditional OPF approach, especially as the system size increases, making the OPF method impractical.
Figure 6: The visualization effectively emphasizes the active power of all scenario samples, both before and after adjustment, across all ten transmission sections of the IEEE 118-bus system using our method.
Figure 6: The visualization effectively emphasizes the active power of all scenario samples, both before and after adjustment, across all ten transmission sections of the IEEE 118-bus system using our method.
Figure 7: The visualization effectively emphasizes the economic cost of all scenario samples, both before and after adjustment, across all ten transmission sections of the IEEE 118-bus system using our method.
4.5 Visualization Analysis
To better understand the superiority of the proposed Powerformer, we present illustrations that depict active power and economic cost across the 10 transmission sections.
Active Power Comparison. By comparing the active power of all scenario samples before and after Powerformer adjustment, across all ten transmission sections on the IEEE 118-bus system, we can find that all test samples are successfully adjusted into the pre-determined range. Among these cases shown in Fig. 7, the majority involve adjusting the situation where the power flow exceeds the limit to align with the range of the upper bound, thereby guaranteeing the safe operation of the power system and ensuring its long-term stability.
Economic Cost Comparison. By comparing the economic cost of all scenario samples before and after Powerformer adjustment on all ten transmission sections of the IEEE 118-bus system, we can find that the economic cost of all test samples has been reduced. As shown in Fig. 7, after undergoing further adjustment through Powerformer, the economic cost of nearly every sample has been effectively reduced, with no instances of cost increase. Hence, it can be inferred that our method reliably aims to reduce economic cost.
Figure 8: The performance of our proposed Powerformer and its ablations on the IEEE 118-bus system.
4.6 Ablation Study
By conducting thorough ablation studies and performing a detailed analysis, we go deeper insights into our proposed Powerformer (PF) architecture. We carefully assess and quantify the contributions of each component, revealing how they individually impact the overall effectiveness and superiority of the algorithm. Fig. 8 presents the comparison results obtained from an array of ablation experiments conducted on IEEE 118-bus system with 10-section task. These research findings provide valuable insights, clarifying the roles of each component in improving performance.
Powerformer-S (PF-S). This method only utilizes the representation of Value component $\boldsymbol{V}$ (Fig. 2) as the final representation of the system state. In contrast to Powerformer-E, it involves the addition of an additional factorization operation. By analyzing the results in Fig. 8, it is evident that comparing Powerformer-S with Powerformer highlights the inadequacy of relying solely on state representations. This emphasizes the importance of incorporating section information.
5 Conclusion
In this paper, we propose and present a specialized Transformer architecture, named Powerformer, which integrates the transmission section power flow information in order to address the issue of power flow adjustment. The utilization of our proposed architecture facilitates controlled power dispatch, leading to improved stability of the power system. Specifically, our proposed Powerformer incorporates transmission section information into the existing Transformer and develops a section-adaptive attention mechanism. Furthermore, our approach disentangles power system graph data into independent electrical factor graphs to conduct a multi-factor attention mechanism, and employs graph neural networks to extract system state features while considering the graph topological structure of the power system. This allows for the robust representations of the sectional power system state, enabling the development of efficient strategies to effectively manage and adjust transmission section power flow. In our future work, we will focus on developing more efficient architectures for factorization and attribution to enhance the performance in adjusting transmission section power flow.

### v4

@@ Appendix C Comparison for both 4-section and 10-section adjustment task
To demonstrate the superiority of our method, we thoroughly analyze and compare our method with various popular approaches in DRL framework, as well as a well-established traditional optimal power flow (OPF) adjustment method. Moreover, in order to thoroughly validate the effectiveness of our Powerformer, we design the additional 4-section power flow adjustment tasks, which aims to deal with the top four ranked sections in terms of transmission capacity among a total of ten sections.
@@
Through analyzing the learning curve in Fig. 9, our proposed method has successfully achieved the best test success rate among all considered architectures on three different systems. We can analyze and compare the superiority of our method based on the difficulty levels of different tasks, especially in the more challenging task of 10-section adjustment (Fig. 9(c)9(d) and 9(e)), where the effectiveness of our method is highlighted. Furthermore, from the perspective of different system cases, our method performs better on large-scale datasets. This is because our method is able to learn better representations of sectional power system states, thus avoiding the coupling of state features. We can comparatively analyze the learning curve on the IEEE 118-bus and European 9241-bus systems, and the significant differences in the effectiveness of our method can be directly observed.
@@
From a quantifiable perspective of comparison adjustment algorithms, we leverage the comparative results of success rates and economic costs pertaining to 4-section and 10-section adjustment tasks, as evidenced in Table 5. Specifically, we further analyze the results by examining different power system cases. For the IEEE 118-bus system, Powerformer outperforms all other methods with a success rate of 99.67%, significantly surpassing Dueling DQN (91.17%) on the 4-section task. As complexity increases in the 10-section task, Powerformer maintains a success rate of 98.19%, outperforming A2C (80.14%) by a substantial margin. For the China 300-bus system, Powerformer continues to exhibit a satisfying level of performance in comparison to the baseline approaches, accomplishing an impressively high test success rate of $99.73\%$ and $97.36\%$ on the 4-section and 10-section tasks, respectively. For the European-9241 system, our method consistently demonstrates exceptional performance, achieving an impressive success rate of 68.97% and 68.25% when considering the test success rates of 4-section and 10-section tasks. In contrast, alternative reinforcement learning methods fail to surpass the 35% threshold for the complex 10-section task. Additionally, our method consistently demonstrates its exceptional ability to achieve relatively low economic costs across all scenarios.
Appendix D More Detailed Analyses
D.1. A Discussion from the Perspective of Computer Science
According to the Fig. 2, our Powerformer adopts a Transformer based architecture to integrate the section information, which is the the main distinction from the existing architectures. This is the first attempt to use the transformer architecture in the field of transmission section power flow adjustment. Specifically, in the existing approaches, the input for state representation learning is homogeneous, while our method can use two heterogeneous electrical information as input. The detailed analyses about section-adaptive attention mechanism, state factorization, GNN integration with transformers, and complexity analysis are as follows:
Section-adaptive Attention Mechanism. Traditional attention-mask techniques primarily focus on preventing the model from accessing future information during training. However, these techniques typically rely on self-attention, which limit their ability to integrate heterogeneous transmission section information. Our section-adaptive approach differs in that it dynamically optimizes attention distribution based on the current section power flow. This allows for a more sophisticated adaptation to the varying demands of the system, enhancing decision-making for resource allocation.
State Factorization. The state factorization and the multi-head (MHA) mechanism are different. The state factors aims to decouple the input into the clear electrical factors (i.e., active power, reactive power, voltage magnitude, and phase angle). They each possess explicit physical meaningful information. This is crucial for understanding the contribution of different factors to the state representation of the power system, which also enhances the interpretability of the achieved representation. As defined in Eq. 11 and Eq. 13, our softmax operates between different factors rather than different nodes. In MHA , the electrical factors are coupling together, making it challenging for each individual attention head to interpret distinct electrical factor significances.
GNN Integration with Transformers. (1) For the task of transmission section power system adjustment, it is often to adjust the generators that are close to the section. Therefore, local information is important for the current section adjustment, and we need to use GNN to capture local information. However, as the power flow is adjusted, the overall operating conditions of the power system changes, and it also affects electrical equipment that is far away. Thus, we need to use Transformer to capture global information. These are essential for developing targeted operational strategies. (2) By leveraging the capability of GNN, our approach dynamically adapts to topological changes such as maintenance, upgrades, or unforeseen emergencies. For example, when a three-phase short circuit occurs on a transmission line, GNN can accurately capture the changes in local information. This helps to maintain robustness against real-world challenges.
Complexity Analysis. The time complexity of Powerformer can be analyzed based on Multi-factor Section-adaptive Attention (MFSA) and Graph Neural Network (GNN) propagation. MFSA is ${O}(nd)$, where $n$ represents the number of bus nodes and $d$ represents the dimensionality of the embeddings. On the other hand, GNN propagation is ${O}(nd^{2}+md)$, where $m$ represents the number of transmission lines. As a result, Powerformer exhibits a time complexity of ${O}(nd+nd^{2}+md)$, while the traditional Transformer is ${O}(n^{2}d)$.
D.2. A Discussion from the Perspective of Power System
In the field of transmission section power flow adjustment, the use of traditional optimal power flow (OPF) methods cannot consider the complexity of real-world scenarios. Additionally, with the expansion of the power system, OPF also raises concerns regarding real-time performance. Therefore, the deep reinforcement learning (DRL) based methods has emerged as an innovative alternative for power dispatch. Specifically, our proposed Powerformer demonstrates its effectiveness not only based on the success rate and economic benefits of the dispatch in three cases but also in terms of its fast adjustment time, which meets the demands of real-world scenarios. Moreover, we provide more discussions and analyses about the application in distribution systems, PV systems, embedded distributed storage systems, and the challenge of data drift.
Distribution Systems. In distribution systems, our model can be conceptualized as ’feeder-adaptive,’ focusing on crucial feeders to control switches that significantly reduce system losses. This adaptation not only serves as load flow adjustment but also demonstrates the model’s applicability in network reconfiguration tasks.
Photovoltaic (PV) Systems. For PV systems, our architecture can help in optimizing the flow of generated solar power, which is highly dependent on variable factors such as sunlight availability. The model’s adaptive attention mechanisms can dynamically adjust to changes in generation capacity, enhancing the efficiency of power distribution and reducing wastage.
Embedded Distributed Storage Systems. In scenarios involving distributed storage systems, our adaptive attention mechanism efficiently identifies and prioritizes the most economically beneficial and supportive energy storage units during complex variations in system operating conditions. By understanding and predicting usage patterns, the system can optimize storage resources to balance the grid during peak and off-peak hours, thereby stabilizing the system and extending battery life.
Data Drift Challenge. In this paper, we focus on the transmission section power flow adjustment, thus not involving data drift. Since data drift is a highly challenging task, which is typically approached as an independent task, involving a detailed analysis and requiring specific designs to effectively address it. Given the established Powerformer framework of this paper, one solution involves a dataset perspective. This includes creating a training dataset that incorporates a variety of special events to ensure that the data covers a diverse range of scenarios. By doing so, the model can learn from these unique situations, thereby enhancing its robustness and ability to handle real-world complexities effectively.
(a) IEEE 118-bus System
(b) Realistic 300-bus System
(c) European 9241-bus System
Figure 10. The performance of our proposed Powerformer and its ablation study on all three power systems.
D.3. A comparison with the existing Graph Transformer
The concept of combining GNNs with Transformers has been previously explored in recent years. In this study, we conducted a comparative analysis by considering three classical methods, i.e, Graph Transformer (Yun et al., 2019), Graphormer (Ying et al., 2021) and GraphFormer (Yang et al., 2021), to highlight the novelty of our work. Graph Transformer (Yun et al., 2019) focuses on node-level prediction within heterogeneous graphs. It utilizes the Transformer’s Query, Key, and Value to generate an adjacency matrix. Subsequently, a GNN employs this matrix to learn node representations. Thus, this method does not actually capture global information, as it uses the Q, K, and V solely to learn an adjacency matrix suitable for the task. Graphormer (Ying et al., 2021) focuses on graph-level prediction by encoding the structural information of the graph into both the input and attention matrices. It then leverages a Transformer model to learn graph representations. As a result, this method does not use GNN within the framework to capture the local graph structure. Instead, it employs the existing techniques such as centrality encoding in input matrix, alongside spatial and edge encoding in attention matrix. GraphFormer (Yang et al., 2021) treats each node in the graph as an independent sample rather than just a token, and then uses a GNN to aggregate the information processed by the Transformer. Thus, this method does not capture the global information of the entire graph using the transformer, but only focuses on the global level between all tokens within the nodes. Next, we provide the comparison analysis from the perspectives of implementation, effectiveness, and application to specific challenges as follows:
Implementation. (1) Graph Transformer employ the Query, Key, and Value to learn adjacency matrix and then use GNN to aggregate the neighborhoods; On the other hand, we employ the Q, K, and V for global information and use GNN to capture the local information. (2) Graphormer embed the graph structure information in the input and attention matrix; On the other hand, we employ GNN as propagation function to capture the local graph information. (3) GraphFormer employ Transformer to capture global level between all tokens within the nodes; On the other hand, we capture global information of the entire graph.
Effectiveness. (1) Graph Transformer is demenstrated in node-level prediction within heterogeneous graphs, which is not for this paper’s graph-level representation. Thus, it is not inappropriate to obtain a comparative analysis about effectiveness. (2) Graphormer encoding structure information into the input and attention matrices, which requires processing and will incur additional O($n^{2}$) computational costs; On the hand, our method employ GNN, which is simple and effective, and only adds linear complexity. Moreover, using GNNs is a better choice. If we were to incorporate structural information into the representations like Graphormer does, it would actually introduce noise, as it would affect the meaning of the original electronic factors in the representations. (3) GraphFormer is proposed for node-level prediction within textual graphs, which is not for this paper’s graph-level representation. Thus, it is not inappropriate to obtain a comparative analysis about effectiveness.
Application to specific challenges. (1) Graph Transformer is designed for heterogeneous, with a specific challenge being the selection of the most important meta-path for GNN. (2) Graphormer faces the challenge of integrating graph structural information into the traditional Transformer architecture. (3) GraphFormer faces the challenge of incorporating textual information within a graph structure. Distinct from these methods, our approach is specifically designed for the power system domain, incorporating transmission section information. The primary challenge here is the effective fusion of section information to maintain accuracy and efficiency. Secondly, we address the complexity of feature entanglement, striving to disentangle into different electronical factors and clarify factor weights for better model performance and insights. Thirdly, to effectively implement dynamic operational strategies in targeted power systems, it is essential to simultaneously utilize both local and global information. This approach allows for precise adjustments to generators close to the section, while also enabling the monitoring and control of electrical equipment located at more distant points within the system.
Appendix E The reason for selecting Dueling DQN
DQN and its variants are off-policy algorithms, while A2C and PPO are on-policy algorithms. Off-policy algorithm enables the agent to update its policy based on the interaction samples from the historical policy, but on-policy algorithm only uses the interaction samples from the current policy. Therefore, the sample efficiency of off-policy algorithm is higher than the on-policy algorithm, which is beneficial in the power system power flow adjustment. In power systems, the environment can be highly stochastic due to unpredictable events such as sudden changes in load or generation. DQN’s off-policy nature allows it to effectively learn from older experiences that may capture a variety of these events, whereas PPO may be more focused on recent experiences that might not be as varied. In the experiments of our original manuscript, the results of 9241-bus System in Table 2 also show that Dueling DQN exhibits superior performances compared to A2C and PPO in the largest power system. In Table 6, we provide experimental comparisons on the IEEE 118 dataset to support our choice of Dueling DQN as the basic backone.
Table 6. The comparison of Dueling DQN and A2C with Powerformer on IEEE 118-bus system and China 300-bus systems.
Method 118-bus System (10-section Task) 300-bus System (10-section Task)
Success Rate (%) Economic Cost (\$) Success Rate (%) Economic Cost (\$)
A2C + Powerformer 90.75 (-7.44) 626,207  (+4,009) 89.21 (-8.15) 972,523  (+2,841)
Dueling DQN + Powerformer 98.19 622,198 97.36 969,682
Appendix F More Ablation Studies on Three Datasets
To further validate the effectiveness of each component proposed, we conducted ablation experiments on two larger datasets, China 300-bus system and European 9241-bus system. As shown in Fig. 10, the ablation results demonstrate that each component within our Powerformer significantly contributes to the overall performance enhancement. This comprehensive analysis highlights the integral role of each element in achieving the final improvements.